# Path Gradients after Flow Matching

### v1

@@ 1 Introduction
Flow Matching, a method for training CNFs, is based on target samples. It is closely related to diffusion model training and has recently gained traction for its simulation-free training and strong empirical performance, both in generative modeling benchmarks (Lipman et al., 2023; Esser et al., 2024; Jin et al., 2024) and in scientific domains (Stark et al., 2024; Jing et al., 2024; Klein & Noé, 2024). Here, we investigate how incorporating Path Gradients, sampled from the target distribution, enhances CNF performance post flow-matching.
Path gradients are low variance gradient estimators, which have strong theoretical guarantees close to the optimum (Roeder et al., 2017) and incorporate gradient information from both, the variational and the target distribution (Vaitl et al., 2024). While they have been adopted for training Normalizing Flows in the field of Lattice Gauge Theory (Bacchio et al., 2023; Kanwar, 2024; Abbott et al., 2023) and also variational inference (Agrawal & Domke, 2024; Andrade, 2023), they remain underused in Quantum Chemistry. In this work, we explore the potential of Path Gradient fine-tuning for Boltzmann Generators, with promising results. This indicates that even though training with Path Gradients is orders of magnitude slower per iteration, its use of additional information and low variance allows it to outperform Flow Matching within the same computational constraints in fine-tuning.
@@ 2 Method
Generative models have recently garnered significant interest for biochemical applications (Noé et al., 2019; Jumper et al., 2021; Abramson et al., 2024). In this work, we focus on the application of enhancing or replacing MD simulations.
@@ 2.1 Boltzmann Generators
[…]
@@
[…]
@@ 2.2 Continuous Normalizing Flows
[…]
@@
Minimizing Equation 7 is equivalent to minimizing the forward KL divergence up to a constant
[…]
@@ 2.3 Flow Matching
$\mathcal{L}_{\text{FM}}$ $(\theta)=\mathbb{E}_{t\sim\mathcal{U}(0,1),x\sim p_{t}(x)}\left[||v_{\theta}(x,t)-u_{t}(x)||^{2}\right]\,.$ (9)
@@ 2.4 Path Gradients
Introduced in the context of variational auto-encoders (Roeder et al., 2017; Tucker et al., 2019), Path Gradient estimators have improved performance, especially applied to normalizing flows (Agrawal et al., 2020; Vaitl et al., 2022a, b; Agrawal & Domke, 2024). In the field of simulating Lattice Field Theories, they have become a go-to tool for training flows (Bacchio et al., 2023; Kanwar, 2024; Abbott et al., 2023). In these applications, the reverse Kullback-Leibler divergence $KL(q_{\theta}|p)$ is minimized by self-sampling, i.e. without existing samples from the target density, but only on samples from $q_{\theta}$ generated while training.
@@
Roeder et al. (2017) observed, that the martingale score term vanishes in expectation, but has a non-zero variance. Vaitl et al. (2022a) showed that its variance is $\frac{1}{N}I_{\theta}$ for a batch of size $N$ and the Fisher Information $I_{\theta}$ of the variational distribution $q_{\theta}$. The low variance close to the optimum allows the gradient estimator to "stick-the-landing" (Roeder et al., 2017), i.e. having zero variance at the optimum, making it an ideal tool for finetuning. Vaitl et al. (2024) showed that Path Gradient estimators incorporate additional information about the derivative of both target and variational distribution, opposed to the reparametrized gradient gradient estimator (Kingma & Welling, 2014; Mohamed et al., 2020). They hypothesized that this leads to less overfitting and a generally better fit to the target density (see Vaitl (2024)). These theoretical aspects make Path Gradients a promising suitor for fine-tuning Boltzmann Generators.
@@
Recently, Path Gradient estimators have been proposed for samples from distributions different from the variational density $q_{\theta}$ (Bauer & Mnih, 2021). By viewing the forward KL$(p_{1}|q_{\theta})$ at $t=1$ as a reverse KL$(p_{0,\theta}|q_{0})$ at $t=0$, we can straightforwardly apply Path Gradients (Vaitl et al., 2024). Here we assume $p_{0,\theta}$ to be a (Continuous) Normalizing Flow from $p_{1}$ as its base density and $T^{-1}_{\theta}$ to be the parametrized diffeomorphism.
@@ 3 Path Gradients and Flow Matching
To build intuition for the behavior of Flow Matching (FM) versus Path Gradient (PG) estimators, we start with a simple 2D Gaussian Mixture Model. We compare three training strategies: 1) FM only, 2) PG only, and 3) a hybrid approach, using FM to quickly approximate the target distribution $p_{1}$, and subsequently applying PG for fine-tuning. Although slower per training step, PG has strong theoretical guarantees near the optimum.
@@
As shown in Figure 1, FM training rapidly improves initially but soon reaches a plateau with slow improvements after. PG training in contrast, progresses more slowly but eventually reaches and surpasses the FM plateau. The hybrid strategy, i.e. beginning with FM and switching to PG, results in the best performance, combining fast convergence with improved final accuracy.
Interestingly, as shown in Figure 2, applying PG after FM has little effect on the MSE loss defined in Equation 10, suggesting that PG fine-tunes the variational distribution $q_{\theta}$ without significantly altering the dynamics $v_{\theta}$. Nonetheless, this subtle adjustment leads to visibly better samples and improved density matching (cf. Figure 1).
@@
In the 2D case, we compute the divergence term in Equation 6 exactly by directly evaluating the trace of the Jacobian. However, this approach scales quadratically with the number of dimensions and quickly becomes computationally prohibitive in higher-dimensional settings. A practical solution is Hutchinson’s trace estimator (Hutchinson, 1989), which provides us faster but noisy estimators for the divergence.
@@
In order to have constant memory irrespective of the number of function evaluations, we adapt the augmented adjoint method (Vaitl et al., 2022b) for Path Gradient estimators for the forward KL divergence Equation 16
@@
as well as the ODEs 5 and 19 per integration step. Even for a single time-step, this is more resource intensive than Flow Matching, which only requires the derivative of the vector field $\frac{\partial v_{\theta}(x_{t},t)}{\partial\theta}$. Thus, although the memory demands do not increase with additional integration steps of the ODE, the memory required for Path Gradients is greater than with Flow Matching. Since the computational load associated with calculating the terms varies across different architectures, we discuss them individually for every experiment.
@@ 4 Experiments
Motivated by the improvements on the toy model, we now turn to more challenging problems, namely the experiments done in Klein et al. (2023b) and Klein & Noé (2024). To this end, we use the same architecture, a CNF with an EGNN (E(n) Equivariant Graph Neural Network) (Satorras et al., 2021; Garcia Satorras et al., 2021) for the vector field. For more details on the model architecture see Section A.2. While we maintain the initial training process with Flow Matching, we change the later training procedure to include Path Gradients. As in Klein et al. (2023b); Klein & Noé (2024) we investigate optimal transport Flow Matching (OT FM), equivariant optimal transport Flow Matching (EQ OT FM), as well as naive Flow Matching (standard FM).
We first evaluate the models on two many particle system with pair-wise Lennard-Jones interactions with 13 and 55 particles. In addition we investigate Alanine dipeptide (AD2) at $T=300K$ both with a classical and a semi-empirical force field (XTB). In contrast to (Klein et al., 2023b; Klein & Noé, 2024), we do not bias the training data towards the positive $\varphi$ state. The bias was originally introduced to make learning problem simpler, as the unlikely state is more prominent (Klein et al., 2023b). This, however, skews the evaluation, since the metrics, namely ESS and Negative log likelihood (NLL), are based on the target distribution, not the biased one. Moreover, this assumes system-specific knowledge of slow-varying degrees of freedom, information that is typically unavailable for unknown systems. For more details on the datasets see Section A.3.
We investigate two scenarios: First, fine-tuning with Path Gradients compared to fine-tuning with OT Flow Matching given the same limited computational resources in Section 4.1. Second, we apply Path Gradients at the end to maximize performance after pre-training with Standard FM, OT FM and EQ OT FM, see Section 4.2. In the second case, we additionally examine the ODE integration length to understand how PG fine-tuning influences the model.
@@ 4.1 Fine-tuning using the same limited resources
We enforce comparable memory and wall-time constraints to those used for Flow Matching alone by adjusting batch size and employing gradient accumulation. Consequently, our fine-tuning runs with Path Gradients complete in less wall-time and occupy a similar memory footprint. For full training statistics and implementation details, see Section A.4 and Section A.5.
Replicating the experiments in Klein et al. (2023b), we evaluate how path-gradient fine-tuning impacts model performance. Table 1 presents a comparison between models trained with pure Flow Matching and those optimized using the hybrid approach of FM pre-training and path-gradient fine-tuning.
@@
[…]
@@
(Klein et al., 2023b) Only FM $-107.89\pm 0.07$ $0.74\pm 0.33$ $0.01\pm 0.01$
Hybrid (ours) $-107.77\pm 0.18$ $0.25\pm 0.37$ $0.01\pm 0.01$
[…]
@@
Figure 3: Reverse ESS, NLL and trajectory length for Flows trained with standard FM, Optimal Transport and Equivariant Transport during fine-tuning with Path Gradients. We can observe that fine-tuning largely leaves the trajectory length unchanged, while substantially improving performance. Mean $\pm$ sterr over three runs.
@@ 4.2 Effect of Path Gradient fine-tuning on flow trajectory length
We use the provided saved models from (Klein et al., 2023b) for LJ13 and investigate the performance and flowed trajectory length during fine-tuning with Path Gradients. We used a batch-size of 256 and trained with Path Gradients for 10 epochs. In Figure 3, we can see that the different pre-trained models have similar NLLs and ESS, but differ in trajectory lengths. Fine-tuning reliably improves the performance of the flows with relative little change to the trajectory lengths. For the full statistics see Section A.7. We observe a similar pattern in the other systems investigated. For alanine dipeptide, for instance, generated samples show significantly improved bond-length and energy distributions, yet the global conformational landscape, captured by the $\varphi$ and $\psi$ dihedral-angle distributions in the Ramachandran plot, remains mostly unchanged (see Figure 4 and Figure 6). These experiments show that our proposed hybrid approach is perfectly suited for maximizing performance while keeping the properties of the model.
@@ 5 Discussion
We have presented a hybrid training strategy for Boltzmann Generators that uses Flow Matching and Path Gradients. We have shown how to efficiently compute Path Gradients with constant memory and without the need for additional samples. While substantially slower per training step, Path Gradients are a powerful tool for fine-tuning, leveraging first-order information from the energy function – an underexplored avenue in the scientific machine learning community. Our results demonstrate that Path Gradients can significantly improve sample quality on the same computational budget as Flow Matching, when the model is already reasonably close to the target distribution. Our experiments show that Path Gradients only apply minor changes to the flow trajectory and to the variational distribution $q_{\theta}$ while still substantially increasing the sampling efficiency and performance. Given our significant improvements, we expect Path Gradients or related methods to become state of the art for Boltzmann Generators and generative models in science.

### v2

@@ 1 Introduction
Flow Matching, a method for training CNFs, is based on target samples. It is closely related to diffusion model training and has recently gained traction for its simulation-free training and strong empirical performance, both in generative modeling benchmarks (Lipman et al., 2023; Esser et al., 2024; Jin et al., 2024) and in scientific domains (Stark et al., 2024; Jing et al., 2024; Klein & Noé, 2024). Here, we investigate how incorporating Path Gradients on samples from the target distribution, enhances CNF performance post flow-matching.
Path gradients are low variance gradient estimators, which have strong theoretical guarantees close to the optimum (Roeder et al., 2017) and incorporate gradient information from both, the variational and the target distribution (Vaitl et al., 2024). While they have been adopted for training Normalizing Flows in the field of Lattice Gauge Theory (Bacchio et al., 2023; Kanwar, 2024; Abbott et al., 2023) and also variational inference (Agrawal & Domke, 2024; Andrade, 2023), they remain underused in Biochemistry. In this work, we explore the potential of Path Gradient fine-tuning for Boltzmann Generators with promising results. This indicates that even though training with Path Gradients is orders of magnitude slower per iteration, its use of additional information and low variance allows it to outperform Flow Matching within the same computational constraints in fine-tuning.
@@ 2 Method
Generative models have recently garnered significant interest for biochemical applications (Noé et al., 2019; Jumper et al., 2021; Abramson et al., 2024). In this work, we focus on the application of enhancing or replacing molecular dynamics simulations.
@@ 2.1 Boltzmann Generators
with an energy $U(x)$ and an unknown normalization factor $Z=\int\exp(-U(x))dx$. The Boltzmann distribution describes the probability of states in a physical system at thermal equilibrium. Traditional sampling methods, such as Markov Chain Monte Carlo (MCMC) and Molecular Dynamics (MD) simulations, generate samples sequentially. In contrast, Boltzmann Generators are designed to produce independent samples directly from the target distribution. There are many different instances of Boltzmann Generators, many focus on molecular systems (Dibak et al., 2022; Köhler et al., 2021; Midgley et al., 2023; Ding & Zhang, 2021b, a; Kim et al., 2024; Rizzi et al., 2023; Tamagnone et al., 2024; Schönle et al., 2024; Klein & Noé, 2024; Tan et al., 2025), while others are trained to sample lattice and other many particle systems (Wirnsberger et al., 2020; Ahmad & Cai, 2022; Nicoli et al., 2020, 2021; Schebek et al., 2024; Abbott et al., 2023; Kanwar, 2024).
@@
$Z=\int\exp(-U(x))dx=\mathbb{E}_{q_{\theta}(x)}\left[\frac{p(x)}{q_{\theta}(x)}\right]$ (3)
@@ 2.2 Continuous Normalizing Flows
$=\log q_{0}(x_{0})-\int_{0}^{1}\text{tr}\left(\frac{\partial v_{\theta}(x_{t},t)}{\partial x_{t}}\right)dt\,.$ (6)
@@
Minimizing Equation˜7 is equivalent to minimizing the forward KL divergence up to a constant
$KL(p|q_{\theta})=\mathbb{E}_{p(x)}\left[\log p(x)-\log q_{\theta}(x)\right]\overset{c}{=}\mathcal{L}_{\text{ML}}(\theta)\,.$ (8)
@@ 2.3 Flow Matching
[…]
@@ 2.4 Path Gradients
Introduced in the context of variational auto-encoders (Roeder et al., 2017; Tucker et al., 2019), Path Gradient estimators have improved performance, especially when applied to normalizing flows (Agrawal et al., 2020; Vaitl et al., 2022a, b; Agrawal & Domke, 2024). In the field of simulating Lattice Field Theories, they have become a go-to tool for training flows (Bacchio et al., 2023; Kanwar, 2024; Abbott et al., 2023). In these applications, the reverse Kullback-Leibler divergence $KL(q_{\theta}|p)$ is minimized by self-sampling, i.e. without existing samples from the target density, but only on samples from $q_{\theta}$ generated while training.
@@
Roeder et al. (2017) observed, that the martingale score term vanishes in expectation, but has a non-zero variance. Vaitl et al. (2022a) showed that its variance is $\frac{1}{N}I_{\theta}$ for a batch of size $N$ and the Fisher Information $I_{\theta}$ of the variational distribution $q_{\theta}$. The low variance close to the optimum allows the gradient estimator to "stick-the-landing" (Roeder et al., 2017), i.e. having zero variance at the optimum, making it an ideal tool for finetuning. Vaitl et al. (2024) showed that Path Gradient estimators incorporate additional information about the derivative of both target and variational distribution, opposed to the reparametrized gradient gradient estimator (Kingma & Welling, 2014; Mohamed et al., 2020) (see Section˜A.11.1 for a detailed explanation). They hypothesized that this leads to less overfitting and a generally better fit to the target density (see Vaitl (2024)). These theoretical aspects make Path Gradients a promising suitor for fine-tuning Boltzmann Generators.
@@
[…]
@@
[…]
Since the publication Vaitl et al. (2022b), Flow Matching has been established as the de facto training method for CNFs. Our work investigates and combines the performance of Flow Matching and Path Gradients. Specifically, we investigate the effect of Path Gradients on the inner workings of the CNF, like e.g. trajectory length, while Vaitl et al. (2022b) simply looked at the performance compared to standard self-sampling losses.
In Section˜A.11, we zoom in on the properties of the different estimators. We show that while Path Gradients exhibit zero variance at the optimum, generically, Flow Matching does not. We further explicate the theoretical properties of the existing estimators for training via Maximum Likelihood, Path Gradients and Flow Matching.
@@ 3 Path Gradients and Flow Matching
To build an intuition for the behavior of Flow Matching (FM) versus Path Gradient (PG) estimators, we start with a simple 2D Gaussian Mixture Model. We compare three training strategies: 1) FM only, 2) PG only, and 3) a hybrid approach, using FM to quickly approximate the target distribution $p_{1}$, and subsequently applying PG for fine-tuning. Although slower per training step, PG has strong theoretical guarantees near the optimum.
@@
As shown in Figure˜1, FM training rapidly improves initially but soon reaches a plateau with slow improvements after. PG training in contrast, progresses more slowly but eventually reaches and surpasses the FM plateau. The hybrid strategy, i.e. beginning with FM and switching to PG, results in the best performance, combining fast convergence with improved final accuracy.
Interestingly, as shown in Figure˜2, applying PG after FM has little effect on the MSE loss defined in Equation˜10, suggesting that PG fine-tunes the variational distribution $q_{\theta}$ without significantly altering the dynamics $v_{\theta}$. Nonetheless, this subtle adjustment leads to visibly better samples and improved density matching (cf. Figure˜1).
@@
In the 2D case, we compute the divergence term in Equation˜6 exactly by directly evaluating the trace of the Jacobian. However, this approach scales quadratically with the number of dimensions and quickly becomes computationally prohibitive in higher-dimensional settings. A practical solution is Hutchinson’s trace estimator (Hutchinson, 1989), which provides us faster but noisy estimators for the divergence.
@@
In order to have constant memory irrespective of the number of function evaluations, we adapt the augmented adjoint method (Vaitl et al., 2022b) for Path Gradient estimators for the forward KL divergence Equation˜16
@@
as well as the ODEs 5 and 19 per integration step. Even for training on a single time-step, this is more resource intensive than Flow Matching, which only requires the derivative of the vector field $\frac{\partial v_{\theta}(x_{t},t)}{\partial\theta}$. Thus, although the memory demands do not increase with additional integration steps of the ODE, the memory required for Path Gradients is greater than with Flow Matching. Since the computational load associated with calculating the terms varies across different architectures, we discuss them individually for every experiment.
@@ 4 Experiments
Motivated by the improvements on the toy model, we now turn to more challenging problems, namely the experiments done in Klein et al. (2023b) and Klein & Noé (2024). To this end, we use the same architecture, a CNF with an EGNN (E(n) Equivariant Graph Neural Network) (Satorras et al., 2021; Garcia Satorras et al., 2021) for the vector field. The architecture in Klein & Noé (2024) is mostly the same as in Klein et al. (2023b), with the main difference of the encoding of the atoms. While the model in Klein et al. (2023b) treats the same atom type as indistinguishable, the model in Klein & Noé (2024) encodes nearly all atoms differently. Only hydrogens bound to the same atom are treated as indistinguishable. For more details on the model architecture see Section˜A.2. While we maintain the architecture and the initial training process with Flow Matching, we change the later training procedure to include Path Gradients. As in Klein et al. (2023b) we investigate optimal transport Flow Matching (OT FM), – and for the second set of experiments on LJ13 – equivariant optimal transport Flow Matching (EQ OT FM), and naive Flow Matching (standard FM).
We first evaluate the models on two many particle system with pair-wise Lennard-Jones interactions with 13 and 55 particles. In addition we investigate Alanine dipeptide (AD2) at $T=300K$ both with a classical and a semi-empirical force field (XTB). In contrast to (Klein et al., 2023b; Klein & Noé, 2024), we do not bias the training data towards the positive $\varphi$ state. The bias was originally introduced to make learning problem simpler, as the unlikely state is more prominent (Klein et al., 2023b). This, however, skews the evaluation, since the metrics, namely ESS and Negative log likelihood (NLL), are based on the target distribution, not the biased one. Moreover, this assumes system-specific knowledge of slow-varying degrees of freedom, information that is typically unavailable for unknown systems. Finally, we also investigate the 2AA dataset consisting of dipeptides simulated at $T=310K$ with a classical force field (Klein et al., 2023a), to evaluate the influence of PG in a transferability. For more details on the datasets see Section˜A.3. We published code for replicating our experiments 222github.com/lenz3000/path-grads-after-fm. In total, we investigate three scenarios:
  • First, fine-tuning with Path Gradients compared to fine-tuning with OT Flow Matching given the same limited computational resources in Section˜4.1.
  • Second, we apply Path Gradients on the FM-trained models using unlimited resources to maximize performance after training with Standard FM, OT FM and EQ OT FM as done in Klein et al. (2023b), see Section˜4.2. In the second case, we additionally examine the ODE integration length to understand how PG fine-tuning influences the model.
  • Finally, we test if the path gradient fine-tuning improves results for transferability in the setting by investigating transferable Boltzmann Generators (Klein & Noé, 2024) trained and evaluated on dipeptides, see Section˜4.3.
@@ 4.1 Fine-tuning using the same limited resources
We enforce comparable memory and wall-time constraints to those used for Flow Matching alone by adjusting batch size and employing gradient accumulation. Consequently, our fine-tuning runs with Path Gradients complete in less wall-time and occupy a similar memory footprint. For full training statistics and implementation details, see Section˜A.4 and Section˜A.5.
Replicating the experiments in Klein et al. (2023b), we evaluate how path-gradient fine-tuning impacts model performance. Table 1 presents a comparison between models trained with pure OT Flow Matching and those optimized using the hybrid approach of OT FM pre-training and path-gradient fine-tuning.
@@
Model Training type NLL $(\downarrow)$ ESSq $\%(\uparrow)$ ESSp $\%(\uparrow)$
@@
(Klein et al., 2023b) Only FM $107.89\pm 0.07$ $74\pm 0.33 0.0$$01\pm 0.01$
Hybrid (ours) $107.77\pm 0.18$ $25\pm 0.37 0.0$$01\pm 0.01$
(Klein & Noé, 2024) Only FM $125.75\pm 0.01$ $3.38\pm 0.50$ $2.65\pm 0.50$
@@
Figure 3: Reverse ESS, NLL and trajectory length for Flows trained with standard FM, Optimal Transport and Equivariant Transport during fine-tuning with Path Gradients on LJ13. We can observe that fine-tuning largely leaves the trajectory length unchanged, while substantially improving performance. Mean $\pm$ sterr over three runs.
@@ 4.2 Effect of Path Gradient fine-tuning on flow trajectory length
We use the provided saved models from (Klein et al., 2023b) for LJ13 and investigate the performance and flowed trajectory length during fine-tuning with Path Gradients. We used a batch-size of 256 and trained with Path Gradients for 10 epochs. In Figure˜3, we can see that the different pre-trained models have similar NLLs and ESS, but differ in trajectory lengths. Fine-tuning reliably improves the performance of the flows with relative little change to the trajectory lengths. For the full statistics see Section˜A.7. We observe a similar pattern in the other systems investigated. For alanine dipeptide, for instance, generated samples show significantly improved bond-length and energy distributions, yet the global conformational landscape, captured by the $\varphi$ and $\psi$ dihedral-angle distributions in the Ramachandran plot, remains mostly unchanged (see Figure˜4 and Figure˜6).
These experiments show that our proposed hybrid approach is perfectly suited for maximizing performance while keeping the properties of the model.
4.3 Transferability on dipeptides
Finally, we also investigate whether fine-tuning with Path Gradients improves performance in the transferable setting. To this end we fine-tune Transferable Boltzmann Generators (TBG) (Klein & Noé, 2024), which were trained on 200 different dipeptides and evaluate on 16 unseen ones, like in Klein & Noé (2024). The experiments show that fine-tuning with path gradients improves the NLL and energies for all evaluated unseen test dipeptides. Further, on average the ESSq is improved by around 23% to an efficiency of $9.79\%$. For more details see Section˜A.10.
@@ 5 Discussion
We have presented a hybrid training strategy for Boltzmann Generators that uses Flow Matching and Path Gradients. We have shown how to efficiently compute Path Gradients with constant memory and without the need for additional samples. While substantially slower per training step, Path Gradients are a powerful tool for fine-tuning, leveraging first-order information from the energy function – an underexplored avenue in the scientific machine learning community. Our results demonstrate that Path Gradients can significantly improve sample quality on the same computational budget as Flow Matching, when the model is already reasonably close to the target distribution. Our experiments show that Path Gradients only apply minor changes to the flow trajectory and to the variational distribution $q_{\theta}$ while still substantially increasing the sampling efficiency and performance.